OpenStudy (anonymous):

The graph passes through (2,4) and (3, 9)

OpenStudy (saifoo.khan):

what u have to find? is it a straight line?

OpenStudy (anonymous):

The linear functions

OpenStudy (anonymous):

Sorry

OpenStudy (saifoo.khan):

do u want the equation?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

the direction just say find the linear function satifyuing the given conditions

OpenStudy (saifoo.khan):

\[ y = 5 x-6\]

OpenStudy (anonymous):

how did you come up with that if I might ask

OpenStudy (saifoo.khan):

Sure. it says to find that function, (2,4) and (3, 9) first of all find it's slope. m = (9-4)/(3-2) m = 5 then y intercept = y = 5x +b 4 = 5(2) + b b = -6 finally, y = mx + b y = 4x -6

OpenStudy (anonymous):

ok what is the formula for the y intercept or is there one?

OpenStudy (saifoo.khan):

No, there is no formula for it. to find the y-intercept, you have to insert the value of slope and any one cordinate. finally, solve for "b: to get the answer.

OpenStudy (anonymous):

ok that makes sence

OpenStudy (anonymous):

Thanks for your help

OpenStudy (saifoo.khan):

No problem! :) can i ask which grade math is that?

OpenStudy (anonymous):

pre calc why?

OpenStudy (saifoo.khan):

ohh..

OpenStudy (anonymous):

yea fun fun

OpenStudy (saifoo.khan):

college?

OpenStudy (saifoo.khan):

pre calc is easy!!

OpenStudy (anonymous):

first year kinda struggling with it. The concepts are hard when it is all online and no teacher instruction at all.

OpenStudy (saifoo.khan):

damn!! u are doing online college? :o then i can understand the hardships!!

OpenStudy (anonymous):

That has been my biggest challenge

OpenStudy (anonymous):

yep not fun but when you live where i do and there is not a college around and you have a family that is what you do.

OpenStudy (anonymous):

I think without this site I would be in a world or hurt about now.

OpenStudy (saifoo.khan):

Yea, correct. i teach many online college people.. and help them with their homeworks.. they have similar similar situation.. and they too dont have good teachers..